OpenStudy (anonymous):

Help please 4x – 3y = 25 –3x + 8y = 10

OpenStudy (anonymous):

x=10 and y=5

OpenStudy (anonymous):

multiply 4x – 3y = 25 by 3 –3x + 8y = 10 by 4 12x – 9y = 75 –12x + 32y = 40 ------------------- 23y = 115 y = 115/23 then 4x = 25 + 3y 4x = 25 + 3 (115/23) 4x = 920/23 x = 230/23 x = 10

OpenStudy (anonymous):

therefore x = 10 and y = 115/23 = 5

OpenStudy (anonymous):

thanks corec you're awesome

OpenStudy (anonymous):

=) Your welcome!